OpenStudy (anonymous):

How do you solve a quadratic formula, I know the formula, but the steps (example; x^2 + x – 4 = 0. )

OpenStudy (rajee_sam):

In the formula you have "a', 'b' and 'c' right.

OpenStudy (anonymous):

yeah

OpenStudy (rajee_sam):

in your example a = 1 (The coefficient of x^2), b = 1(The coefficient of x) and c = -4 (The Constant Term). So plug in the values in the formula and find your x.

OpenStudy (rajee_sam):

There will two values for X

OpenStudy (rajee_sam):

\[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

I know the part to get some finished, like plugging it in, but once i multiply and get to the (2a) what do I do? what do i do with the (-b)

OpenStudy (jhannybean):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]\[\frac{ -1 \pm \sqrt{(1)^2-4(1)(-4)} }{ 2(1) }\] I'll let you go on from there.

OpenStudy (anonymous):

yeah for that it would be (1) - (-4) (-4) all over (2)

OpenStudy (jhannybean):

mmhmm, and you'd get two answers as your solution

OpenStudy (anonymous):

which basically you just multiply the -4's and get -16? then divide it by 2? which is -4? is that all?

OpenStudy (rajee_sam):

sorry that does not look right @1010Dunno1010

OpenStudy (rajee_sam):

inside the sqrt. it is 1- 4(1)(-4) = 1 + 16 = 17

OpenStudy (jhannybean):

\[\frac{ -1-\sqrt{17} }{ 2 }\] and \[\frac{ -1+\sqrt{17} }{ 2}\]

OpenStudy (jhannybean):

Thank you rajee :)

OpenStudy (anonymous):

Oh I see!